OpenStudy (anonymous):

A building has a ramp to its front doors to accommodate the handicapped. The angle of elevation of the ramp must be no greater than 4º. If the height from the ground to the front doors is 5 feet, what is the minimum length of the ramp? (Round to the nearest tenth.)

OpenStudy (xkehaulanix):

tan(4º)=5/x xtan(4º)=5 x=5/tan(4º) \[x \approx 71.5 \]